OpenStudy (anonymous):

Compare and contrast, in complete sentences, the difference between the products of x^3x^4 and (x^3)^4. Provide the expanded form and product of each.

OpenStudy (anonymous):

(x^3)(x^4)=x^(3+4)=x^7, (x^3)^4=x^((3)(4))=x^12

OpenStudy (anonymous):

That's not a complete sentence, but WHATEVA I DO WHAT I WANT.

OpenStudy (anonymous):

lol! cartmen quote :D And thank you